# Data Driven Decision Making with Time Series and Spatio-temporal Data

### v1

@@ II-A Data Foundations
Definition 4 (Image Sequence).
@@ II-B Data Governance
Uncertainty Quantification: This process aims to capture uncertainty related to time and space using probability distributions. Histograms or Gaussian mixture models are often used to represent distributions, as they are able to approximate distributions without assumptions on the type of distribution. We cover a dynamic, uncertainty graph model in the context of intelligent transportation. Traffic flow is inherently uncertain and dynamic, influenced by factors like accidents and traffic volume. This can be modeled using $(I,D)$ pairs, denoting that the travel speed follows distribution $D$ within time interval $I$. We discuss two paradigms: the edge-centric paradigm [15] and path-centric paradigm [4]. The former assigns distributions to edges, treating the edge distributions as independent, while the latter captures the distribution correlations along paths, balancing efficiency and precision. Uncertainty quantification enhances data reliability, providing a robust foundation for further analytics and decision-making, e.g., decision making under uncertainty [6].
Multi-modal Fusion: Time series and spatio-temporal data exhibit multi-source, heterogeneous, multi-modal traits, with significant correlations across time and space. Data fusion across modalities improves data quality by consolidating data, laying a robust foundation for further analysis and decision-making. Recent multi-modal fusion approaches can be summarized into four streams: feature-based, alignment-based, contrast-based, and generation-based methods [16]. A prime instance of alignment-based fusion is Map Matching [17], which aligns vehicle trajectory data with corresponding map data. This process can eliminate trajectory noise as well as compensate for gaps in trajectories, thereby enhancing data quality and analytical and decision-making processes. Another example is feature-based fusion of multi-modal data, such as historical traffic, weather data, and point of interest, for traffic forecasting [18, 19]. Advances in representation learning and foundational models enable the capture of multi-modal spatio-temporal data as vectors or embeddings [20, 21, 22]. For example, MM-Path enhances path representations by aligning embeddings of the road network and satellite image of paths [23]. By constructing temporal and spatial indices, we can improve the efficiency of analytics on diverse multi-modal spatio-temporal datasets.
@@ II-C Data Analytics
Automation: Traditional time series and spatio-temporal analysis methods are often crafted by human experts, which is resource-intensive. Automated methods offer a practical alternative: a search space is created for a specific task, upon which algorithms can identify and assemble automatically the most fitting models for the specific task and dataset. For instance, automated correlated time series forecasting automates the design of network architectures and hyperparameter selection [24, 25, 26, 27, 28]. Such automation not only enhances accuracy across diverse datasets and forecasting settings but also facilitates the discovery of optimal models that adhere to additional constraints, e.g., model sizes.
@@
Generality: Methods using deep learning are capable of superior performance in time-series and spatio-temporal analytics, but they depend on high-quality representations. Representation learning, which requires extensive labeled data, poses challenges in data acquisition and does not generalize across tasks [29]. We consider general representation learning frameworks, pre-trained on abundant unlabeled data, which can be fine-tuned with minimal labeled data for different spatio-temporal tasks. These include unsupervised contrastive learning [30], weakly-supervised contrastive learning [31], masked autoencoders with relational inference [32], and cross-modal contrastive learning [23]. These frameworks demonstrate generality and reduce the dependence on large amounts of labeled data. With the emergence of Artificial General Intelligence, we also consider how to leverage Large Language Models (LLMs) to enhance the zero-shot adaptability of deep learning models for time series forecasting [20, 33] and spatio-temporal data mining [21, 22].
@@
Robustness: Robustness in time-series and spatio-temporal analysis encompasses handling anomalous data, data imbalance, distribution shifts, and multi-scale adaptation. Traditional unsupervised anomaly detection algorithms assume implicitly that training occurs on fully-clean data, which is rarely available in practice. We thus cover robust anomaly detection [34, 35] that operates effectively on noisy training data. Considering data imbalance, we cover an adversarial domain adaptation method [36] that models unbalanced data patterns despite data size discrepancies. Addressing distribution shifts, e.g., due to changes in the physical world such as new roads, we cover continuous learning [37, 38] tailored to spatio-temporal data dynamics. In addition, we cover means to model heterogeneous temporal patterns caused by temporal distribution shifts, e.g., in streaming data or non-stationary time series [39]. Lastly, we cover multi-scale modeling using pathways [40] and ensemble learning strategies [41, 42] that enhance robustness by adaptively selecting and combining multiple scales for analytics.
@@
Explainability: Effective decision-making requires both accuracy and explainability in data analytics. Although deep learning exhibits state-of-the-art accuracy at time series and spatio-temporal analyses, its “black box” nature hampers explainability. Considering this, we focus on two important aspects: how to quantify the explainability of different methods and how to improve the explainability of deep learning methods. We present a posthoc explainability metric [35] for the assessment of autoencoder-based anomaly detection models. Furthermore, we consider how to leverage neural networks for feature extraction and integrate extracted features with interpretable models to augment overall prediction clarity [43]. Additionally, we consider how to enhance explainability by tracking temporal associations among time series [44, 45] and employing causal models to predict future correlations [46]. Lastly, we cover how LLMs offer remarkable explainability in the realm of spatio-temporal data mining [21, 20], enabling the injection of domain knowledge in language forms and facilitating the understanding of patterns within time series and spatio-temporal data.
@@
Resource efficiency: Due to increasing privacy concerns, analytics are increasingly performed on personal edge devices rather than having to upload personal data to potentially untrustworthy servers. Contemporary analysis methods prioritize accuracy, resulting in the reliance on computationally intensive model architectures that are unsuitable for resource-constrained edge devices. Such methods also do not align with green computing principles. To cover this aspect, we present LightPath [32], a knowledge distillation-based framework for efficient spatio-temporal data representation, and LightTS [47], an automated, lightweight temporal analysis framework capable of adapting quantization levels to align with memory limitations of edge devices. QCore enables continuous calibrations for quantized models, addressing both resource efficiency and robustness on distribution shifts [48]. TimeDC studies a time series dataset condensation method that is able to compress large time series into a smaller counterpart while maintaining key properties [49]. This facilitates data analytics in resource constrained environments.
@@ II-D Data-Driven Decision Making
Learning-based Decision Making: These strategies simulate the decision patterns of experts, enabling non-experts to replicate expert-level decisions [56]. For instance, professional taxi drivers possess an intimate understanding of urban traffic, making their routing decisions valuable knowledge. By analyzing the trajectories of expert drivers, it is possible to enable human drivers and autonomous vehicles to mimic their behavior. Beyond simple imitation, this strategy involves dissecting and enhancing the determinants of expert decisions to achieve superior outcomes.
@@ II-E Research Directions
Most existing studies are highly specialized, focusing on individual cases, specific locations and time. This specificity leads to models with limited applicability, high usage costs, and reduced scalability. To broaden the technology’s reach and reduce barriers to its use, the future of temporal spatio-temporal decision intelligence may include: (1) Pre-training and few-shot learning: Leveraging pre-training on diverse tasks and extensive datasets across various spatio-temporal contexts can yield foundational models with broad applicability. Such models may facilitate few-shot or zero-shot learning when subjected to novel tasks and environments. (2) Generative model potential: Advances in generative models may improve performance in temporal and spatio-temporal decision-making. The strong generalization capabilities and precision in data generation of such models may lead to better spatio-temporal transferability. (3) Integration with LLMs: Merging temporal and spatio-temporal decision-making with LLMs may enable more intuitive and seamless interactions through natural language.
@@ III Presenters
Yuxuan Liang is a tenure-track Assistant Professor at Hong Kong University of Science and Technology (Guangzhou), working on spatio-temporal data mining and urban computing. He has published 70+ papers in prestigious venues (e.g., TPAMI, AI, TKDE, KDD, NeurIPS, and WWW). He served as organizer or co-chair of Workshop on Urban Computing (Urbcomp@KDD) and Workshop on AI for Time Series (AI4TS@IJCAI).

### v2

@@ II-A Data Foundations
Definition 4 (Image Sequences).
@@ II-B Data Governance
Uncertainty Quantification: This process aims to capture uncertainty related to time and space using probability distributions. Histograms or Gaussian mixture models are often used to represent distributions due to their capacity to approximate distributions without assumptions on the type of distribution. We cover a dynamic, uncertainty graph model in the context of intelligent transportation. Traffic flow is inherently uncertain and dynamic, influenced by factors such as accidents and traffic volume. This can be modeled using $(I,D)$ pairs, where travel speed follows distribution $D$ over time interval $I$. We discuss two paradigms: the edge-centric paradigm [15] and path-centric paradigm [4]. The former assigns distributions to edges, treating the edge distributions as independent, while the latter captures the distribution correlations along paths. Uncertainty quantification enhances data reliability, providing a robust foundation for further analytics and decision-making, e.g., decision making under uncertainty [6].
Multi-modal Fusion: Time series and spatio-temporal data exhibit multi-source, heterogeneous, multi-modal characteristics, with significant correlations across time and space. Data fusion across modalities improves data quality by consolidating data, laying a robust foundation for further analysis and decision-making. Recent multi-modal fusion approaches can be summarized into four streams: feature-based, alignment-based, contrast-based, and generation-based methods [16]. A prime instance of alignment-based fusion is Map Matching [17], which aligns vehicle trajectory data with corresponding map data. This process can eliminate trajectory noise as well as compensate for gaps in trajectories, thereby enhancing data quality and analytical and decision-making processes. Another example is feature-based fusion of multi-modal data, such as historical traffic, weather data, and point of interest, for traffic forecasting [18, 19]. Advances in representation learning and foundational models enable the capture of multi-modal spatio-temporal data as vectors or embeddings [20, 21, 22]. For example, MM-Path enhances path representations by aligning embeddings of the road network and satellite image of paths [23]. By constructing temporal and spatial indices, we can improve the efficiency of analytics on diverse multi-modal spatio-temporal datasets.
@@ II-C Data Analytics
Automation: Traditional time series and spatio-temporal analysis methods are often crafted by human experts, which is resource-intensive. Automated methods offer a practical alternative: a search space is created for a specific task, upon which algorithms can identify and automatically assemble the most fitting models for the specific task and dataset. For instance, automated correlated time series forecasting automates the design of network architectures and hyperparameter selection [24, 25, 26, 27, 28]. Such automation not only enhances accuracy across diverse datasets and forecasting settings but also facilitates the discovery of optimal models that adhere to additional constraints, e.g., model sizes.
@@
Generality: Methods using deep learning are capable of superior performance in time-series and spatio-temporal analytics, but they depend on high-quality representations. Representation learning, which requires extensive labeled data, poses challenges in data acquisition and does not generalize across tasks [29]. We consider general representation learning frameworks, pre-trained on abundant unlabeled data, that can be fine-tuned with limited labeled data for different spatio-temporal tasks. These include unsupervised contrastive learning [30], weakly-supervised contrastive learning [31], masked autoencoders with relational inference [32], and cross-modal contrastive learning [23]. For time series tasks, we propose general models for forecasting [33] and anomaly detection [34] to eliminate the need for target-domain-specific training data. These frameworks demonstrate generality and reduce the dependence on large amounts of labeled data. With the emergence of Artificial General Intelligence, we also consider how to leverage Large Language Models (LLMs) to enhance the zero-shot adaptability for time series forecasting [20, 35] and spatio-temporal data mining [21, 22].
@@
Robustness: Robustness in time-series and spatio-temporal analysis encompasses handling anomalous data, data imbalance, distribution shifts, and multi-scale adaptation. Traditional unsupervised anomaly detection algorithms assume implicitly that training occurs on fully-clean data, which is rarely available in practice. We thus cover robust anomaly detection [36, 37, 38, 39] that operates effectively on noisy training data. Considering data imbalance, we cover an adversarial domain adaptation method [40] that models unbalanced data patterns despite data size discrepancies. Addressing distribution shifts, e.g., due to changes in the physical world such as new roads, we cover continuous learning [41, 42] tailored to spatio-temporal data dynamics. In addition, we cover means to model heterogeneous temporal patterns caused by temporal distribution shifts, e.g., in streaming data or non-stationary time series [43]. Lastly, we cover multi-scale modeling using pathways [44] and ensemble learning strategies [45, 46] that enhance robustness by adaptively selecting and combining multiple scales for analytics.
@@
Explainability: Effective decision-making requires both accuracy and explainability in data analytics. Although deep learning exhibits state-of-the-art accuracy at time series and spatio-temporal analyses, its “black box” nature hampers explainability. Considering this, we focus on two important aspects: how to quantify the explainability of different methods and how to improve the explainability of deep learning methods. We present a posthoc explainability metric [37] for the assessment of autoencoder-based anomaly detection models. Furthermore, we consider how to leverage neural networks for feature extraction and integrate extracted features with interpretable models to augment overall prediction clarity [47]. Additionally, we consider how to improve the explainability by tracking temporal associations among time series [48, 49] and employing causal models to predict future correlations [50]. We integrate physical equations into data-driven models to improve the explainability of air quality prediction [51]. Lastly, we cover how LLMs offer remarkable explainability in the realm of spatio-temporal data mining [21, 20], enabling the injection of domain knowledge into language forms and facilitating the understanding of patterns within time series and spatio-temporal data.
@@
Resource efficiency: Due to increasing privacy concerns, analytics are increasingly performed on personal edge devices rather than having to upload personal data to potentially untrustworthy servers. Contemporary analysis methods prioritize accuracy, resulting in the reliance on computationally intensive model architectures that are unsuitable for resource-constrained edge devices. Such methods also do not align with green computing principles. To cover this aspect, we present LightPath [32], a knowledge distillation-based framework for efficient spatio-temporal data representation, and LightTS [52], an automated, lightweight temporal analysis framework capable of adapting quantization levels to align with memory limitations of edge devices. QCore [53] enables continuous calibrations for quantized models, addressing both resource efficiency and robustness on distribution shifts. EnhancingDFQ [54] improves the accuracy of data-free model quantization by increasing sample-data diversity. TimeDC studies a time series dataset condensation method that is able to compress large time series into a smaller counterpart while maintaining key properties [55]. This facilitates data analytics in resource constrained environments.
@@ II-D Data-Driven Decision Making
Learning-based Decision Making: These strategies simulate the decision patterns of experts, enabling non-experts to replicate expert-level decisions. For instance, we mimic expert taxi drivers’ routing strategies by extracting valuable knowledge from their historical trajectories and decision patterns [62]. HiSSD [63] incorporates cooperation temporal learning within a hierarchical architecture, identifying shared and specific skills across tasks, thereby enhancing policy generalization.
@@ II-E Research Directions
Most existing studies are highly specialized, focusing on individual cases, specific locations, and time. This specificity leads to models with limited applicability, high costs, and reduced scalability. To broaden the technology’s reach and reduce barriers to its use, the future of temporal and spatio-temporal decision intelligence may include: (1) Pre-training and few-shot learning: Leveraging pre-training on diverse tasks and extensive datasets across various spatio-temporal contexts can yield foundational models with broad applicability. Such models may facilitate few-shot or zero-shot learning when applied to novel tasks and domains. (2) Generative model potential: Advances in generative models may improve performance in temporal and spatio-temporal decision-making. The strong generalization capabilities and precision in data generation of such models could enable better spatio-temporal transferability. (3) Integration with LLMs: Merging temporal and spatio-temporal decision-making with LLMs may enable more intuitive and seamless interactions through natural language.
@@ III Presenters
Yuxuan Liang is a tenure-track Assistant Professor at Hong Kong University of Science and Technology (Guangzhou), working on spatio-temporal data mining and urban computing. He has published 70+ papers in prestigious venues (e.g., TPAMI, AI, TKDE, KDD, NeurIPS, and WWW).